### v1

# From Laboratory to Real-World Applications: Benchmarking Agentic Code Reasoning at the Repository Level

Abstract
As large language models (LLMs) evolve into autonomous agents, evaluating Repository-Level Reasoning, the ability to maintain logical consistency across massive, interdependent file systems, has become critical. Current benchmarks typically fluctuate between isolated code snippets and black-box evaluations. We present RepoReason, a white-box diagnostic benchmark centered on Abductive Assertion Verification. To eliminate memorization while preserving authentic logical depth, we implement an Execution-Driven Mutation framework that utilizes the environment as a Semantic Oracle to regenerate ground-truth states. Furthermore, we establish a fine-grained diagnostic system using dynamic program slicing, quantifying reasoning via three orthogonal metrics: ESV (Reading Load), MCL (Simulation Depth), and DFI (Integration Width). Comprehensive evaluations of frontier models (e.g., Claude-4.5-Sonnet, DeepSeek-v3.1-Terminus) reveal a prevalent Aggregation Deficit, where integration width serves as the primary cognitive bottleneck. Our findings provide granular white-box insights for optimizing the next generation of agentic software engineering.
From Laboratory to Real-World Applications: Benchmarking Agentic Code Reasoning at the Repository Level
Jia Li The Chinese University of Hong Kong Hong Kong, China linsayli@link.cuhk.edu.hk          Yuxin Su Sun Yat-sen University Zhuhai, China suyx35@mail.sysu.edu.cn
Michael R. Lyu The Chinese University of Hong Kong Hong Kong, China lyu@cse.cuhk.edu.hk
1 Introduction
The paradigm of software engineering is undergoing a fundamental shift: from human-centric coding to agentic autonomous development (Jiang et al., 2025; Hou et al., 2024). As Large Language Models (LLMs) evolve into autonomous software agents, the core competency required has escalated from performing localized execution reasoning within isolated code snippets—akin to controlled laboratory exercises—to maintaining logical consistency across extensive, interdependent file systems—a capability we define as Repository-Level Reasoning. This capability imposes profound challenges on LLMs, demanding not only code synthesis but also the agency to explore vast contexts to pinpoint target definitions, aggregate multi-source information across complex dependencies, and perform long-chain reasoning to accurately track and compute state mutations.
However, a critical gap exists between this growing demand and current evaluation methods. Existing benchmarks largely fall into two extremes. On one side, function-level reasoning datasets, such as CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024), study code execution reasoning within isolated snippets. While these provide analytical rigor, they operate in a laboratory setting that lacks the cross-file dependencies and complex architectural contexts found in real-world software engineering. On the other side, frameworks like SWE-bench (Jimenez et al., 2024) evaluate agents in wild repository environments but function as black-boxes: they check whether a problem is solved but offer little insight into why an agent fails. Without detailed diagnosis, the path to improving agent reasoning remains unclear.
To address this gap, we introduce RepoReason, a repository-level code reasoning benchmark designed as a white-box diagnostic tool. Unlike traditional generation tasks, RepoReason shifts to a verification-centric approach: “Given the complex execution history of this repository, what is the current system state?” By requiring models to derive deterministic values that satisfy assertions rather than writing implementation logic, we separate core logical reasoning from syntactic noise (e.g., formatting errors, indentation slips, or API name hallucinations). This ensures that every failure reflects a genuine deficit in reasoning rather than superficial coding errors.
To ensure the benchmark reflects real-world complexity, we extract task instances from a curated set of mainstream Python repositories (e.g., toolz (Developers, 2025b), sympy (Team, 2025c), jinja2 (Pallets, 2025)), challenging models with authentic, deep logic. Beyond static code extraction, we perform concrete execution of the repositories’ built-in unit test suites to capture granular runtime traces. By designing structural metrics based on these traces—such as stack depth, function call density, and cross-file dependency breadth—we quantitatively filter and retain only the most complex reasoning tasks.
Constructing such a benchmark requires navigating between two distinct failure modes of current evaluations: Spurious Correlation (guessing via shallow patterns) and Rote Memorization (recalling via pre-training data). To address this, we propose a two-stage defense mechanism. First, to prevent shortcut learning via visual pattern matching, we reject raw trace values and instead identify Unit Test Assertions as our semantic anchors. While raw traces are often localized and easily inferred from immediate code patterns, assertions function as semantic milestones that encapsulate high-level developer intent, strategically positioned to validate the system state after complex logical operations. They force Abductive Reasoning (Josephson and Josephson, 1994), requiring the model to reconstruct the preceding execution history across complex call stacks to deduce the verified state.
However, a significant obstacle remains: many assertions within these mainstream repositories’ unit tests have already been encountered by LLMs during pre-training. This leads to a risk of data leakage (Deng et al., 2024). To resolve this, we implement an Execution-Driven Mutation methodology. We treat the code execution environment as a Semantic Oracle. By keeping the reasoning logic (the call graph filtered via structural metrics) intact but perturbing the program inputs, we effectively sever the model’s memory retrieval path. We then re-execute the entire repository context to capture the new ground-truth state, subsequently masking these verified values to create cloze-style reasoning tasks. Finally, to ensure these captured states are objectively evaluable, we introduce a Deterministic Value Protocol, which filters runtime data to eliminate representational drift and ensure stable, unique ground-truth values. This design ensures that while the reasoning logic remains authentic and complex, the specific state values are unmemorized and deterministic.
Finally, through an extensive evaluation of a diverse range of frontier models including Claude-Sonnet-4.5 and DeepSeek-v3.1-Terminus, we demonstrate that even top-tier agents face severe performance degradation when confronted with high-entropy logic. Our results unveil a critical Aggregation Deficit where accuracy drops sharply as integration width increases, as well as a Cliff Effect in reading comprehension beyond 600 lines and a significant loss of consistency beyond 100 execution steps. These insights highlight that next-generation agents must transcend local reasoning and focus on high-breadth information synthesis and long-chain state consistency.
In summary, this paper makes the following contributions:
  • From Prediction to Abductive Verification: We pioneer a repository-level reasoning paradigm that pivots from direct outcome prediction to abductive assertion verification. By centering tasks on masked assertions within interdependent file systems, we successfully decouple core architectural intelligence from surface-level syntactic noise.
  • The Semantic Oracle Framework: We resolve the Richness-Contamination Dilemma through an Execution-Driven Mutation framework. By utilizing the execution environment as a Semantic Oracle, we sever data leakage paths via input perturbation and real-time state regeneration, ensuring tasks remain logically authentic yet unmemorized.
  • White-box Diagnostic Instrument: Beyond end-to-end scoring, we introduce a granular diagnostic system powered by dynamic program slicing. We formalize three orthogonal cognitive metrics (Reading Load, ESV; Simulation Depth, MCL; Integration Width, DFI) that map reasoning failures to specific cognitive bottlenecks: context overload, state tracking deficits, and aggregation barriers.
  • Empirical Unveiling of Aggregation Deficits: Through extensive evaluation of frontier models, we provide the first quantitative analysis of agentic reasoning trajectories across massive codebases (up to 776k+ LoC). Our findings identify DFI as the primary bottleneck for next-generation agents, empirical confirming a severe Aggregation Deficit in information synthesis.
Table 1: Comparison of RepoReason with existing code reasoning and understanding benchmarks.
Benchmark Repo-Lvl Scale (LoC) Evaluation Paradigm Metric Granularity Diagnostic Depth
HumanEval (Chen et al., 2021) $\times$ $\sim$10 Code Generation Pass/Fail Black-box
MBPP (Austin et al., 2021) $\times$ $\sim$10 Code Generation Pass/Fail Black-box
ClassEval (Du et al., 2023) $\times$ $\sim$40 Code Generation Pass/Fail Black-box
CRUXEval (Gu et al., 2024) $\times$ $\sim$10 Execution Prediction Pass/Fail Black-box
REval (Chen et al., 2024) $\times$ $\sim$10–40 Runtime Behavior Pass/Fail White-box (Local)
CORE (Xie et al., 2025) $\times$ 21–100 Static Analysis Pass/Fail White-box (Static)
SWE-bench (Jimenez et al., 2024) $\checkmark$ 1k – 1M+ Issue Resolution Pass/Fail Black-box
RepoReason (Ours) $\checkmark$ 1.2k – 775k+ Assertion Verification Cognitive Metrics White-box (Global)
2 Related Work
2.1 Code Generation
Existing code benchmarks have evolved from function-level synthesis (e.g., HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021)) to class-level and repository-level completion (e.g., ClassEval (Du et al., 2023), RepoBench (Liu et al., 2024)). While these frameworks increasingly reflect real-world complexity, they primarily target syntactic completion—the ability to produce correct code text. In contrast, RepoReason shifts the focus to semantic reasoning. By requiring the deduction of deterministic system states via assertion verification, it isolates core logical reasoning from the syntactic noise and hallucination risks inherent in code generation.
2.2 Code Reasoning
Recent benchmarks like CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) pioneered execution reasoning but are largely confined to isolated, “lab-scale” code snippets. While tools like CORE (Xie et al., 2025) incorporate static analysis tasks to evaluate reasoning, they lack the structural depth and code breadth (e.g., deep inheritance, cross-module imports) characteristic of large-scale projects. RepoReason bridges this gap by applying deep execution simulation to massive repository scopes, challenging models to maintain logical consistency across extensive file systems.
2.3 Agentic Software Engineering
At the highest level, SWE-bench (Jimenez et al., 2024) evaluates agents on resolving GitHub issues. However, as an end-to-end evaluation, it acts as a “black-box” indicator of success or failure. RepoReason complements this by targeting the cognitive root causes of failure. By decoupling reasoning from code editing and utilizing fine-grained metrics (ESV, MCL, DFI), it provides the “white-box” diagnostic insights required to pinpoint an agent’s specific bottlenecks in context handling or multi-source integration.
2.4 Summary Comparison
Table 1 provides a multi-dimensional comparison between RepoReason and mainstream code benchmarks, highlighting its unique advantages in diagnostic depth and structural complexity.
Figure 1: The overall architecture and multi-stage pipeline of RepoReason, encompassing repository curation, structural filtering, execution-driven mutation, task instantiation, and diagnostic evaluation.
3 RepoReason: Design and Construction
In this section, we delineate the methodological framework of RepoReason, conceptualizing the construction pipeline as a rigorous five-phase defense system. Our objective is to transform authentic software logic into “white-box” diagnostic tasks while systematically neutralizing systemic vulnerabilities: spurious correlation, rote memorization, and logical superficiality. As illustrated in Figure 1, the pipeline orchestrates sequential gates: (1) Repository Curation to ensure environmental fidelity; (2) Structural Filtering via runtime traces to guarantee logical depth; (3) Execution-Driven Mutation to eliminate parametric leakage; (4) Task Instantiation governed by a deterministic value protocol for stability; and (5) Diagnostic Evaluation establishing fine-grained cognitive profiles. This integration ensures that each task instance necessitates authentic repository-level reasoning and robust state tracking, effectively precluding resolution via shallow pattern matching.
3.1 Design Philosophy: Assertions as Semantic Anchors
We explicitly reject two prevailing paradigms: traditional Input/Output (I/O) Prediction and Raw Trace Querying.
I/O Prediction Inefficiency: The traditional I/O Prediction paradigm (Gu et al., 2024; Chen et al., 2024), while effective for isolated algorithmic puzzles, is fundamentally ill-suited for repository-level reasoning. In authentic software ecosystems, runtime data structures, such as SymPy’s symbolic expression trees or NetworkX’s graph topologies, are often voluminous and complex, lacking the clear, concise boundaries found in synthetic tasks. Consequently, verifying the entire output state is inherently inefficient and noise-laden, as a significant portion of the data consists of dynamic metadata or irrelevant state rather than the target logic.
Spurious Correlation in Raw Traces: A straightforward approach might involve randomly querying variable values at specific locations (e.g., “What is the value of variable $x$ at line 10?”). However, this “Raw Trace Querying” is fundamentally flawed for evaluating deep reasoning. It suffers from shortcut learning, as models can often retrieve answers through shallow static analysis of nearby assignment statements, bypassing the need for genuine execution simulation.
The Solution: Assertions as Semantic Anchors. To overcome these pitfalls, we identify Unit Test Assertions as the optimal semantic units for instantiation. Assertions are not random data points; they are crystallized checkpoints of developer intent, strategically placed to verify critical state transitions after complex operations. By targeting assertions, we filter out implementation noise and focus solely on the logical invariants of the system. We employ a Mask-and-Reason strategy: we take a valid assertion (e.g., assert len(cache) == 5) and mask the ground truth value (e.g., assert len(cache) == <mask>). This transforms the task into Abductive Reasoning: the model cannot find the answer by reading the immediate text. Instead, it must mentally reconstruct the preceding execution history, tracing data flows across files and simulating state mutations, to deduce the only logical value that satisfies the verification condition.
3.2 Foundation: Real-world Repository Curation
To ensure that our “Semantic Anchors” are grounded in reality, we curated a set of mature, actively maintained Python repositories. This stage ensures the fidelity of our benchmark, providing the complex, cross-file dependency structures that synthetic datasets lack.
3.2.1 Repository Selection
We selected mature, actively maintained, and predominantly pure Python open-source projects from PyPI and GitHub. Our selection specifically targets libraries where the core logic is implemented in Python, enabling complete line-level tracing via sys.settrace. Libraries with substantial C/C++/Cython extensions (e.g., NumPy, Pandas, PyTorch, scikit-learn) were explicitly excluded, as their core computational routines remain opaque to Python-level instrumentation.
The selected repositories cover diverse reasoning domains and exhibit diverse complexity, with codebases ranging from approximately 1,200 to over 775,000 lines of code (LoC). This deliberate variation allows for a comprehensive evaluation of LLMs’ code reasoning capabilities across different scales and perspectives. Detailed description and statistics are shown in Table 5, Section A.
3.2.2 File-Level Semantic Filtering
Test selection prioritizes cases requiring multi-step semantic reasoning over superficial pattern matching. We manually review source code at the file level to screen all test modules. This initial stage aims to filter out low-value assertions and ensure Deterministic Verifiability. We systematically excluded test files falling into three specific categories: (1) Trivial API Validation, comprising simple attribute access (e.g., assert url.scheme == ’http’), direct dictionary operations without cross-module dependencies (e.g., assert d.get(’key’) == ’value’), and metadata introspection lacking computational reasoning (e.g., assert num_required_args(lambda x: None) == 1); (2) Shallow Computation Patterns, including direct formula application or lookup tables (e.g., assert quote(’a b’) == ’a%20b’), simple string manipulation without state complexity, and local transformations requiring no cross-module reasoning (e.g., assert x + 0 == x); and (3) Non-deterministic or External Dependencies, such as tests relying on randomness (e.g., assert random.choice(seq) in seq) or external timing or system state (e.g., assert time.time() > start).
3.3 Countering Memorization: The Execution-Driven Mutation Engine
The second line of defense addresses Rote Memorization. Since powerful LLMs have likely memorized public codebases, using raw GitHub code leads to data contamination. Our Execution-Driven Mutation framework serves as an "Anti-Contamination Engine," creating parallel reasoning realities that preserve logic but alter facts.
Unified Semantic and Logic Mutation. We employ a unified mutation strategy utilizing a teacher LLM. In a single step, the model performs both visual changes (e.g., renaming variables, restructuring code, and removing comments) and logic changes (e.g., modifying constants, input data, and parameters) (Jia and Harman, 2010). This combined approach ensures the mutated code is correct and readable, while effectively removing visual clues and invalidating memorized answers. Crucially, it strictly preserves the original API call sequence to the target library to maintain invariant reasoning complexity.
Execution-Driven Ground Truth Regeneration. To ensure the correctness of mutated tests and avoid LLMs hallucinations, we designed an Execution-Driven Assertion Reconstruction framework. First, we use Probe Injection, where the LLMs converts original assertions into print statements formatted as DEBUG_RESULT, turning the test into code that outputs its internal state. Next, we run this code in a controlled environment to capture the actual Runtime Values produced by the logic changes. These values serve as the ground truth, derived directly from code execution rather than model guessing. Finally, we perform Style-Consistent Assertion Generation. Using the captured values and the original code’s style as guides, the LLMs reconstructs the assertions. It mimics the original writing patterns, for example, keeping is for object identity checks instead of changing to ==, or using specific constructors like FiniteSet(1, 2) instead of native sets like {1, 2}. This ensures the new ground truth is mathematically accurate and matches the original codebase’s style.
Strict Validation Gate. We enforce a strict validation process. A mutated test is considered valid only if it meets three criteria: (1) Executability, where the code runs without errors; (2) Assertion Validity, ensuring new assertions pass against the mutated logic; and (3) API Preservation, where the API call sequence matches the original test. Any variant failing these checks is discarded, ensuring all tasks in RepoReason are solvable, deterministic, and maintain original reasoning depth.
3.4 Countering Superficiality: Trace-Based Structural Filtering
The third defense targets Logical Superficiality. Even in complex repositories, many tests remain shallow. To ensure RepoReason evaluates genuine Repository-Level capabilities, we employ Trace-Based Structural Filtering (Phase 2) to identify tasks with sufficient logical depth.
Trace Collection. With the mutated test cases ready, we capture their runtime behavior using a lightweight tracing system based on Python’s built-in sys.settrace() function, which enables line-level execution tracking. Our system records the complete sequence of function calls during the execution of the mutated tests, capturing granular details including function names, module paths, line numbers, call order, call stack depth, function source code, input arguments, return values, and executed lines within each function. Critically, to eliminate noise, we automatically identify the test function as the entry point and focus strictly on the target library’s execution path. Furthermore, to prevent infinite recursion and ensure data manageability, we enforce a maximum call depth of 3, where the test function itself is defined as depth 0.
Function-Level Structural Filtering and Complexity Grading. Instead of traditional hard thresholds, we implemented a weighted scoring system based on captured traces to identify high-density test cases. We calculated a complexity score for each test function:
$\text{Score}=$ $\left(\frac{N_{\text{files}}}{4}\times 0.1\right)+\left(\frac{N_{\text{funcs}}}{15}\times 0.2\right)$ (1)
$+\left(\frac{N_{\text{calls}}}{30}\times 0.5\right)+\left(\frac{D}{4}\times 0.2\right)$
This formula prioritizes computational intensity (Call Count, 50%) and execution depth (Stack Depth, 20%), while balancing coverage breadth (Function Count, 20%) and cross-module dependencies (File Count, 10%). The baselines ($N_{\text{files}}=4,N_{\text{funcs}}=15,N_{\text{calls}}=30,D=4$) are calibrated to the empirical distribution of the target repositories. We primarily select tests rated 3 stars or higher ($\text{Score}\geq 0.30$), filtering out approximately 13% of trivial functions.
3.5 Countering Ambiguity: The Deterministic Value Protocol
To ensure reproducibility and eliminate evaluation noise, we introduce the Deterministic Value Protocol. This protocol enforces that ground-truth answers are unique and stably representable through two programmatic filtering funnels (covering $\sim$90% of samples):
Semantic Determinism Funnel: Automatically filters samples with inherent uncertainty, including: (1) representational drift (e.g., strings containing random memory addresses); (2) external perturbations (e.g., non-deterministic signals like timestamps or UUIDs); and (3) multi-solution semantics (e.g., range comparisons like assert x > 10).
Morphological Determinism Funnel: Mandates that masked values represent concrete runtime states rather than arbitrary code text. A whitelist priority is established: Literals >> Global Constants >> Parameter-Resolvable Constructors. As a special case, sympy allows complex mathematical expressions (e.g., Rational(33, 2) - Rational(11, 2)*sqrt(3)) as valid answers, as the expression itself serves as the standard storage form. Pure variable references or dynamic function calls are strictly rejected.
For descriptive variables used by developers (e.g., assert result == expected), the system activates an LLM Variable Resolver to unfold abstract references into runtime-verified literal forms, ensuring evaluation is anchored to deterministic system states.
3.6 Cognitive Diagnostic Framework
To quantify the intrinsic reasoning complexity of a task, we perform dynamic program slicing (Agrawal and Horgan, 1990) rooted at the assertion to obtain the “Minimal Causal Computational Subgraph” along the actual execution path. Dynamic cross-function backward slicing is obtained by tracing back from the assertion along data and control dependencies within a recorded execution trace to isolate the minimal set of statements that causally influence the observed result. Based on this subgraph, we characterize reasoning intensity using three mutually orthogonal metrics rooted in cognitive load theory (Sweller, 1988).
1. Reading Load (ESV — Effective Sliced Volume): Measures the causally relevant source code volume required to process. We quantify the source volume of computational units (functions/methods) in the slice rather than simple slice line counts, as reasoning requires restoring semantic context (parameters, constraints, preconditions) within stable boundaries. Failures at high ESV signify Comprehension Failure rooted in Context Overload, where the agent loses semantic focus across the extended causal chain.
2. Simulation Depth (MCL — Mutation Chain Length): Measures the number of dynamic state update steps required to reach the target state. We treat each causally relevant statement in the dynamic slice as a state advancement atomic step and sum them weighted by execution frequency. This metric reflects the difficulty of maintaining a virtual machine state. Significant performance deterioration at high MCL indicates Simulation Failure stemming from a State Tracking Deficit, where the agent struggles to accurately synchronize variable states through complex sequences, leading to the breakdown of deep simulation.
3. Integration Width (DFI — Dependency Fan-in): Measures the number of independent upstream logical inputs that collectively determine the critical value at an assertion. It quantifies logical inputs consumed by the execution subgraph but generated outside of it, while explicitly filtering out structural noise such as language keywords or module imports. Performance degradation at high DFI reveals an Integration Failure caused by an Aggregation Deficit in synthesizing disparate logical sources into a coherent conclusion.
Figure 2: Performance trajectories of frontier LLM agents across three orthogonal cognitive metrics: (a) Effective Sliced Volume (ESV), (b) Mutation Chain Length (MCL), and (c) Dependency Fan-in (DFI).
4 Experimental Evaluation
To validate the diagnostic effectiveness of RepoReason, we conducted a comprehensive evaluation using state-of-the-art LLMs.
4.1 Experimental Setup
Agent Framework. We utilized OpenHands (version 0.60.0) (Team, 2025a) as our evaluation platform. Specifically, we configured the ReadOnlyAgent for all tasks. This agent configuration is restricted to file system exploration and code reading operations, preventing any modification to the repository. This setup perfectly aligns with the verification nature of RepoReason, isolating the model’s ability to “read, navigate, and reason” from its ability to “edit or plan.”
Model Selection. We evaluated a diverse set of frontier models representing the latest advancements in code reasoning: the Claude Family (claude-sonnet-4.5 (Anthropic, 2025)), the GPT Family (gpt-5.2 (OpenAI, 2025)), the DeepSeek Family (deepseek-3.1-terminus (DeepSeek-AI, 2025)), the Kimi Family (Kimi-K2 (AI, 2025)), and the Qwen Family (qwen3-coder-480b-a35b (Team, 2025b)).
Inference Configuration. To ensure reproducibility and eliminate generation randomness, we set the temperature to 0 for all model evaluations. Additionally, to accommodate the potentially lengthy reasoning chains required for repository-level tasks, we set the maximum generation length (max_tokens) to 8192 for each inference step.
4.2 Overall Performance: Accuracy across Repositories and Difficulty Levels
We report the Pass@1 accuracy of the four model families across the seven selected repositories. To control for task variance, results are stratified into Easy ($N=1009$), Medium ($N=838$), and Hard ($N=645$) groups based on the LLM’s Self-Perceived Difficulty, with a total sample size of Overall ($N=2492$).
Table 2 summarizes the overall performance of the agents. All models exhibit a clear performance degradation from Easy to Hard tasks, with Claude-Sonnet-4.5 maintaining the highest accuracy across all difficulty levels.
Table 2: Overall Agent Accuracy by Difficulty (%)
Model Overall Easy Medium Hard
Claude-Sonnet-4.5 66.98 80.27 66.79 47.39
DeepSeek-v3.1-Terminus 60.96 74.03 60.02 41.71
GPT-5.2 56.86 75.92 53.94 30.85
Kimi-K2 54.74 69.38 51.79 35.66
Qwen3-Coder-480B 50.56 61.35 49.16 35.50
Through a detailed analysis of Table 3, we observe three distinct categories:
1. Explicit Algorithmic Patterns (Cachetools and Toolz): In repositories such as cachetools and toolz, models demonstrate robust and consistent performance, with mean accuracies ranging from 78.00% to 82.00%. These libraries primarily rely on standard algorithmic paradigms—such as cache eviction and functional data flows—characterized by high logical locality and linear execution transitions. The results indicate that LLMs have effectively mastered these procedural patterns, which currently define the “Cognitive Comfort Zone” for agentic reasoning.
2. Implicit Metaprogramming Logic (Attrs and Jinja2): attrs and jinja2 exhibit a significant performance stratification. Taking attrs as an example, which relies heavily on decorators for implicit class generation (e.g., the dynamic injection of __init__ methods), the performance gap between Claude (72.41%) and Kimi-K2 (43.10%) reveals a fundamental divergence in the internalization of the Python dynamic object model. Frontier models demonstrate the capacity to model runtime behaviors that remain “invisible” as explicit code lines in static text, whereas other models show a marked deficiency when logic is generated programmatically rather than defined literally.
3. High-Entropy Symbolic and Structural Reasoning (SymPy and NetworkX): sympy and networkx represent the primary “Reasoning Ceiling” for current agents. Symbolic transformations in sympy involve abstract, non-linear logical transitions that frequently cause reasoning chains to deviate. Concurrently, networkx imposes a high demand on Structured Working Memory to maintain complex graph topologies. Claude’s superior performance in networkx (79.77%) confirms its advanced capability to maintain accurate “state snapshots” during complex traversals without experiencing disorientation.
Table 3: Agent Accuracy Breakdown by Repository (%)
Model cache (193) toolz (320) yarl (192) netx (404) jinja (325) attrs (58) symp (1000)
Claude-4.5 79.79 82.81 78.65 79.77 67.46 72.41 51.10
Deep-3.1 78.24 77.19 70.83 63.37 74.15 62.07 45.20
GPT-5.2 78.76 72.19 81.77 65.10 66.46 65.52 36.00
Kimi-K2 78.76 62.81 52.08 70.54 60.92 43.10 39.60
Qwen3-Coder 70.98 62.81 59.90 57.43 55.38 43.10 37.00
4.3 Diagnostic Analysis via Cognitive Metrics
To provide a systematic white-box diagnosis of agentic code reasoning, we analyze model performance across three orthogonal cognitive dimensions: Reading Load (ESV), Simulation Depth (MCL), and Integration Width (DFI).
Figure 2 visualizes the performance trajectories of all model families as a function of the three cognitive intensity metrics. We observe a consistent and pronounced inverse relationship across all axes: accuracy systematically degrades as ESV, MCL, and DFI increase. This trend empirically validates our diagnostic framework, demonstrating that RepoReason effectively quantifies the escalation of cognitive load in repository-level reasoning and serves as a robust discriminator for architectural intelligence.
A comparative analysis of the decline slopes identifies DFI as the primary cognitive bottleneck for next-generation software agents. The slope of performance degradation is steepest along the DFI axis across all models. When DFI exceeds 20 independent upstream sources, accuracy for most models, with the exception of Claude, falls below the 40% threshold, revealing an Aggregation Deficit: agents struggle to parallelly hold and synthesize constraints from multiple disparate logical sources.
Regarding context processing, we identify a clear “Cliff Effect” in ESV. Performance remains relatively stable in lower context ranges but exhibits a sharp decline once the volume of causally relevant computational units exceeds 600 LoC. This threshold marks the transition into Context Overload, where agents fail to maintain semantic focus across the sliced causal chain. Similarly, for MCL, we observe a significant deterioration in logical consistency beyond 100 execution steps.
Table 4 provides the precise Pearson correlation coefficients for these trends. The statistical evidence confirms that while reading load and update depth are significant factors, DFI exhibits the strongest negative correlation with accuracy, reaching -0.234 for GPT-5.2. This confirms that Integration Failure is the most severe obstacle in repository-level reasoning, necessitating that future research prioritize enhancing agents’ capabilities in high-breadth information synthesis.
Table 4: Correlation Coefficients between Cognitive Metrics and Accuracy
Model ESV MCL DFI
GPT-5.2 -0.188 -0.158 -0.234
Claude-Sonnet-4.5 -0.161 -0.122 -0.225
DeepSeek-v3.1-Terminus -0.152 -0.122 -0.157
Kimi-K2 -0.130 -0.117 -0.195
Qwen3-Coder-480B -0.148 -0.154 -0.196
Analysis of Table 4 reveals distinct cognitive profiles for each model family, highlighting unique architectural trade-offs. DeepSeek-v3.1 emerges as the most architecturally balanced model, showing exceptional robustness against high Integration Width (DFI, 0.157-0.157), whereas Claude-4.5 excels in simulation depth but remains fragile in multi-source integration (0.225-0.225). Furthermore, while Qwen3-Coder demonstrates superior resilience to context volume (ESV, 0.148-0.148) and performs optimally in handling large-scale contexts, its state-tracking consistency lags behind the leaders, and GPT-5.2 remains globally the most vulnerable model across all dimensions.
5 Limitations
Despite the rigorous design of RepoReason, we identify several areas for future refinement:
Language Scope: Currently, RepoReason focuses primarily on Python. Although Python remains the de facto standard for AI-native development and LLM research, extending this benchmark to languages such as Java, C++, and Rust is a priority for future work. We address this potential limitation by ensuring that our system architecture is built on standard profiling interfaces, which are conceptually portable to other language ecosystems, thereby enabling our "white-box" diagnostic paradigm to cover a broader range of programming environments.
Logical Depth Relative to Base Tests: The reasoning depth of the generated tasks is partially limited by the complexity of the original repository’s test suites. We ensure foundation quality by curating mature projects with exceptionally high test coverage and complexity. The mutation engine then evolves these human-written "fixed" tests into vast permutations of reasoning scenarios, significantly enhancing task diversity.
6 Ethics Statement
RepoReason is constructed entirely from public repositories with open-source licenses that permit research and software usage. During data collection and evaluation, we do not collect metadata regarding repository pull request authors or individual contributors, and our pipeline utilizes only publicly available information accessible via standard APIs. Our work did not involve human subject participation; we did not crowdsource or recruit human workers for any stage of the benchmark construction. All manual validation of task instances, environment setup, and metric calibration were conducted solely by the authors.
Upon publication, we plan to open-source the RepoReason task instances, the Execution-Driven Mutation framework, and the comprehensive model reasoning trajectories to facilitate transparency and community research. We will document each component according to best practices and establish communication channels to engage the community in refining and improving the benchmark.
Appendix A Repository Selection and Statistical Overview
Table 5 presents the detailed statistics and core reasoning scenarios for our selected repositories, where lines of code (LoC) are measured using the cloc tool and specifically represent source code by excluding configuration files, documentation, and metadata.
Table 5: Statistics and Key Reasoning Scenarios of Selected Repositories
Repository LoC Tests Core Reasoning Complexity
cachetools Kemmer (2025) 1.2k 88 Cache eviction (LRU/LFU/TTL), state management
yarl Svetlov (2025) 2.3k 369 URL parsing, normalization, codec consistency
toolz Developers (2025b) 3.7k 92 Func. flow, lazy evaluation, high-order logic
attrs Schlawack (2025) 6.3k 209 Metaprogramming, class gen, attribute validation
jinja2 Pallets (2025) 14k 527 Template compilation, lexical parsing, context flow
networkx Developers (2025a) 67k+ 534 Graph traversal, centrality, complex connectivity
sympy Team (2025c) 776k 913 Symbolic math, solver logic, recursive calculus
Appendix B Other Statistics from Experiment and Dataset
B.1 Diagnostic Insights: Reasoning Scope
Figure 3: Performance of LLM agents across different reasoning scopes.
Figure 3 illustrates the impact of reasoning scope on agent performance. We observe a consistent performance decay as the task transitions from Intra-file Isolation to Repository-Wide Synthesis. This downward trend quantifies the Retrieval Failure described in our diagnostic framework, highlighting a significant Spatial Deficit in current agents’ ability to locate and navigate long-distance dependencies across the repository architecture. Notably, cross-file reasoning tasks constitute the vast majority of RepoReason, further emphasizing its authentic repository-level nature.
B.2 Dataset Characteristics: Cognitive Metrics Distribution
Figure 4: Empirical distribution of the three cognitive metrics across the RepoReason dataset: (a) ESV, (b) MCL, and (c) DFI.
Figure 4 illustrates the empirical distribution of the three cognitive metrics across the RepoReason dataset. All metrics exhibit a characteristic long-tail distribution, indicating a wide range of task difficulties. The Effective Sliced Volume (ESV) distribution ($\mu=393.6,\sigma=706.3$) shows that a significant portion of tasks require navigating hundreds or thousands of lines of code. The Mutation Chain Length (MCL) distribution ($\mu=93.9,\sigma=198.0$) highlights the depth of state tracking required, while the Dependency Fan-in (DFI) distribution ($\mu=8.3,\sigma=7.5$) captures the integration complexity of multi-source information. This diverse distribution ensures that RepoReason provides sufficient data points across various complexity levels for fine-grained diagnostic analysis.
Appendix C Technical Specification of the Agentic Search Framework
This section details the implementation of the Agentic Search evaluation framework, focusing on the specialized prompt engineering for the ReadOnlyAgent and the robust evaluation mechanisms used to verify repository-level reasoning.
C.1 Agentic Search Prompting Strategy
The framework utilizes a template-driven approach to construct tasks for the Agent. As illustrated in Box 1, the prompt is structured to provide necessary execution context while enforcing strict output constraints.
Agentic Search Task Prompt Template Task: Fill in the blank in the Python test code. Blank Assertion: [Blanked_Assertion_Statement] Hint: [Type_Specific_Hint] (e.g., "The answer MUST be a string literal" or "MUST be a constructor call from the source code"). Instructions: The ’___’ in the assertion represents a missing expression (variable name, function call, or literal). You are a ReadOnlyAgent. You can read files and search the code in the repository to infer what this expression should be. Note: Test function code is not visible in the repository, but you can explore the source code functions called by the test and reason to complete the fill-in-the-blank task. Context: - Test function source code: [Source_Code_with_Question_Marker] - Input parameters: [JSON_Serialized_Inputs] Note: The missing expression is likely related to the input parameters or variables defined in the test or the code being tested. Goal: Find the exact expression that should replace ’___’. When you are confident, use the finish tool with the answer. IMPORTANT: The finish tool’s content MUST contain ONLY the answer expression itself, nothing else. Examples: If the answer is a variable: finish with content "result" If the answer is a function call: finish with content "len(items)" If the answer is a string: finish with content ’"hello"’ If the answer is an attribute access: finish with content "C.__attrs__attrs__[1].validator" Do NOT include explanations, reasoning, or any other text in the finish tool content. Only the answer expression.
C.2 Evaluation and Comparison Engine
Given the syntactic flexibility of Python, the framework employs a multi-criteria comparison engine to verify the Agent’s output against ground truth.
C.2.1 Mathematical Equivalence ($\mathcal{M}$)
For tasks involving symbolic expressions, the engine utilizes the SymPy library to verify mathematical identity. Correctness is defined as:
$\text{Correct}\iff$ $\text{Simplify}(P-G)=0$ (2)
$\lor\text{Expand}(P)=\text{Expand}(G)$
where $P$ is the Agent’s prediction and $G$ is the ground truth. This multi-path verification prevents false negatives caused by structural variations.
C.2.2 Structural Normalization ($\mathcal{S}$)
The engine applies recursive normalization to handle implementation-specific data structures:
  • Wrapper Stripping: Automatically unwrap custom container types (e.g., D({...}) $\to$ {...}).
  • Collection Alignment: Standardize whitespace and element ordering in complex literals.
  • Object Mapping: Map specialized collections like defaultdict to standard dictionary equivalents.
C.2.3 Heuristic Semantic Matching
To bridge the gap between precise syntax and reasoning intent, the following heuristics are applied:
  • Constructor Awareness: Recognizes class instantiations required by the logic (e.g., ClassName(args)).
  • Path Qualification: Standardizes module paths, accepting partially qualified references if unambiguous.
  • Quote Invariance: Neutralizes differences in quote types (single, double, or triple).
C.3 Execution Management
The Agent is executed in a sandboxed environment with a strict 1800-second timeout. All tool invocations are captured in a serialized JSON sequence for auditability. An incremental checkpointing mechanism is implemented to ensure data persistence during large-scale evaluations.

### v2

@@
 2.2 Code Reasoning
-Recent benchmarks like CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) pioneered execution reasoning but are largely confined to isolated, “lab-scale” code snippets.
+Recent benchmarks like CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) pioneered execution reasoning but are largely confined to isolated, lab-scale code snippets.
 While tools like CORE (Xie et al., 2025) incorporate static analysis tasks to evaluate reasoning, they lack the structural depth and code breadth (e.g., deep inheritance, cross-module imports) characteristic of large-scale projects.
@@
 At the highest level, SWE-bench (Jimenez et al., 2024) evaluates agents on resolving GitHub issues.
-However, as an end-to-end evaluation, it acts as a “black-box” indicator of success or failure.
+However, as an end-to-end evaluation, it acts as a black-box indicator of success or failure.
 RepoReason complements this by targeting the cognitive root causes of failure.
-By decoupling reasoning from code editing and utilizing fine-grained metrics (ESV, MCL, DFI), it provides the “white-box” diagnostic insights required to pinpoint an agent’s specific bottlenecks in context handling or multi-source integration.
+By decoupling reasoning from code editing and utilizing fine-grained metrics (ESV, MCL, DFI), it provides the white-box diagnostic insights required to pinpoint an agent’s specific bottlenecks in context handling or multi-source integration.
 2.4 Summary Comparison
@@
 In this section, we delineate the methodological framework of RepoReason, conceptualizing the construction pipeline as a rigorous five-phase defense system.
-Our objective is to transform authentic software logic into “white-box” diagnostic tasks while systematically neutralizing systemic vulnerabilities: spurious correlation, rote memorization, and logical superficiality.
+Our objective is to transform authentic software logic into white-box diagnostic tasks while systematically neutralizing systemic vulnerabilities: spurious correlation, rote memorization, and logical superficiality.
 As illustrated in Figure 1, the pipeline orchestrates sequential gates: (1) Repository Curation to ensure environmental fidelity; (2) Structural Filtering via runtime traces to guarantee logical depth; (3) Execution-Driven Mutation to eliminate parametric leakage; (4) Task Instantiation governed by a deterministic value protocol for stability; and (5) Diagnostic Evaluation establishing fine-grained cognitive profiles.
@@
 Spurious Correlation in Raw Traces: A straightforward approach might involve randomly querying variable values at specific locations (e.g., “What is the value of variable $x$ at line 10?”).
-However, this “Raw Trace Querying” is fundamentally flawed for evaluating deep reasoning.
+However, this paradigm of raw trace querying is fundamentally flawed for evaluating deep reasoning.
 It suffers from shortcut learning, as models can often retrieve answers through shallow static analysis of nearby assignment statements, bypassing the need for genuine execution simulation.
@@
 3.2 Foundation: Real-world Repository Curation
-To ensure that our “Semantic Anchors” are grounded in reality, we curated a set of mature, actively maintained Python repositories.
+To ensure that our semantic anchors are grounded in reality, we curated a set of mature, actively maintained Python repositories.
 This stage ensures the fidelity of our benchmark, providing the complex, cross-file dependency structures that synthetic datasets lack.
@@
 3.6 Cognitive Diagnostic Framework
-To quantify the intrinsic reasoning complexity of a task, we perform dynamic program slicing (Agrawal and Horgan, 1990) rooted at the assertion to obtain the “Minimal Causal Computational Subgraph” along the actual execution path.
+To quantify the intrinsic reasoning complexity of a task, we perform dynamic program slicing (Agrawal and Horgan, 1990) rooted at the assertion to obtain the minimal causal computational subgraph along the actual execution path.
 Dynamic cross-function backward slicing is obtained by tracing back from the assertion along data and control dependencies within a recorded execution trace to isolate the minimal set of statements that causally influence the observed result.
@@
 These libraries primarily rely on standard algorithmic paradigms—such as cache eviction and functional data flows—characterized by high logical locality and linear execution transitions.
-The results indicate that LLMs have effectively mastered these procedural patterns, which currently define the “Cognitive Comfort Zone” for agentic reasoning.
+The results indicate that LLMs have effectively mastered these procedural patterns, which currently define the cognitive comfort zone for agentic reasoning.
 2.
@@
 Taking attrs as an example, which relies heavily on decorators for implicit class generation (e.g., the dynamic injection of __init__ methods), the performance gap between Claude (72.41%) and Kimi-K2 (43.10%) reveals a fundamental divergence in the internalization of the Python dynamic object model.
-Frontier models demonstrate the capacity to model runtime behaviors that remain “invisible” as explicit code lines in static text, whereas other models show a marked deficiency when logic is generated programmatically rather than defined literally.
+Frontier models demonstrate the capacity to model runtime behaviors that remain invisible as explicit code lines in static text, whereas other models show a marked deficiency when logic is generated programmatically rather than defined literally.
 3.
-High-Entropy Symbolic and Structural Reasoning (SymPy and NetworkX): sympy and networkx represent the primary “Reasoning Ceiling” for current agents.
+High-Entropy Symbolic and Structural Reasoning (SymPy and NetworkX): sympy and networkx represent the primary reasoning ceiling for current agents.
 Symbolic transformations in sympy involve abstract, non-linear logical transitions that frequently cause reasoning chains to deviate.
-Concurrently, networkx imposes a high demand on Structured Working Memory to maintain complex graph topologies.
-Claude’s superior performance in networkx (79.77%) confirms its advanced capability to maintain accurate “state snapshots” during complex traversals without experiencing disorientation.
+At the same time, networkx requires high structured working memory to track complex graph topologies.
+Claude’s superior performance in networkx (79.77%) confirms its advanced ability to maintain accurate state snapshots during complex traversals without experiencing memory decay or structural confusion.
 Table 3: Agent Accuracy Breakdown by Repository (%)
@@
 Only the answer expression.
+C.1.1 Concrete Example: Literal Integer Answer
+We present a real prompt instance from the cachetools repository to illustrate the complete Agentic Search prompt structure.
+This example demonstrates a literal integer answer, where the Agent must infer the value 2 by reasoning about cache operations.
+Real Agentic Search Prompt Example Task: Fill in the blank in the Python test code.
+Blank Assertion: assert ___ == len(cache) Hint: CRITICAL REQUIREMENT: The answer MUST be a literal integer value.
+IMPORTANT: Return ONLY the literal value, NOT an expression or variable name.
+Type-specific examples: Boolean: True or False (NOT ’x == y’ or ’bool(x)’) String: "hello" (NOT ’str_var’ or ’x.name’) Number: 42 (NOT ’num_var’ or ’len(x)’) List: [1, 2, 3] (NOT ’list_var’ or ’list(x)’) Dict: {"a": 1} (NOT ’dict_var’ or ’dict(x)’) Instructions: The ’___’ in the assertion represents a missing expression (variable name, function call, or literal).
+You are a ReadOnlyAgent.
+You can read files and search the code in the repository to infer what this expression should be.
+Note: Test function code is not visible in the repository, but you can explore the source code functions called by the test and reason to complete the fill-in-the-blank task.
+Context: - Test function source code: class LRUCacheTest(unittest.TestCase, CacheTestMixin): Cache = LRUCache def test_update(self): cache = self.Cache(maxsize=2) cache.update({1: 1, 2: 2}) assert ___ == len(cache) # Question assert ___ == cache[1] assert ___ == cache[2] cache.update({1: 1, 2: 2}) assert ___ == len(cache) assert ___ == cache[1] assert ___ == cache[2] cache.update({1: "a", 2: "b"}) assert ___ == len(cache) assert ___ == cache[1] assert ___ == cache[2] - Input parameters: {"self": "<LRUCacheTest>"} Note: The missing expression is likely related to the input parameters or variables defined in the test or the code being tested.
+Goal: Find the exact expression that should replace ’___’.
+When you are confident, use the finish tool with the answer.
+IMPORTANT: The finish tool’s content MUST contain ONLY the answer expression itself, nothing else.
+Examples: If the answer is a variable: finish with content "result" If the answer is a function call: finish with content "len(items)" If the answer is a string: finish with content ’"hello"’ If the answer is an attribute access: finish with content "C.__attrs__attrs__[1].validator" Do NOT include explanations, reasoning, or any other text in the finish tool content.
+Only the answer expression.
+Analysis. In this example, the Agent must:
+  1. 1.
+    Understand that cache.update({1: 1, 2: 2}) adds two key-value pairs to a cache with maxsize=2
+  2. 2.
+    Reason that after the update, len(cache) equals 2
+  3. 3.
+    Infer that the blank should be filled with the literal integer 2
+  4. 4.
+    Use the finish tool with content "2" (without quotes in the tool content, as it’s a numeric literal)
+This example demonstrates how the prompt guides the Agent to reason about runtime behavior while enforcing strict output constraints for evaluation.
 C.2 Evaluation and Comparison Engine